here is the word problem:
The American society of heating, refrigeration and air conditioning requires 4 complete air changes per hour in public buildings. If a classroom is 30 ft by 35 ft with a 10 ft ceiling, how many cubic feet of air must be moved every hour in this room?
Im at a loss on how to set this question up. Thank you!
Dawn

Answer by rothauserc(4718)

Volume = length * width * height
Volume = 30 * 35 * 10 = 10500 cubic ft
Four complete air changes per hour
10500 * 4 = 42000 cubic feet per hour